InfoAssume thatit takes the US 150 units of labor to produce a table and 30 units of labor to produce a chair Assume it takes the Netherlands 100 units of labor to produce a table and 10 units of labor to produce a chair that each country has 4800 units of labor and before trade, each splits their labor evenly between chair and table production Assume that when the countries decide to trade, the country specializing in production specializes in table production only. that when the countries decide to trade, the country that in chair production uses 1000 units of labor for table production the rest (3800 units) for chair production. Assume that when the countries trade, there are a total of 15 tables traded at a rate of 7 chairs per table As -US SPECIALIZES IN TABLE PRODUCTION NETHERLANDS SPECIALIZES IN CHAIR PRODUCTION QUESTIONS 1. After Trade, How many tables Does the US consume? 2. After Trade, How many chairs does the US consume? 3. After Trade, how many tables does the Netherlands Consume? 4. After Trade, How many chairs does the Netherlands Consume? Trade chair Wethesland S Chair
Answer #1

Answer:

1):We calculate each of the opportunity cost and we come up as follows:

opportunity cost of producing 1 chair in Netherlands = No of tables produced/ No of chairs produced = 24/240 = 0.1
opportunity cost of producing 1 table in Netherlands = No of chairs produced/ No of tables produced = 240/24 = 10
opportunity cost of producing 1 chair in US = No of tables produced/ No of chairs produced = 16/80 = 0.2
opportunity cost of producing 1 table in US = No of chairs produced/ No of tables produced = 80/16 = 5
Hence, US should produce tables and Netherlands should produce chairs because of their respective low opportunity cost. Now, US uses all its labor in producing tables, so it produces 0 chairs and 4800/150 = 32 tables. Ans since Netherlands specializes in producing chairs it produces 3800/10 = 380 chairs and 1000/100 = 10 tables.

After trade US trades 15 tables hence, it consumes 32-15 = 17 tables

2). Since, US gets 7 chairs for each table traded, hence US gets 15*7 = 105 chairs and since it was not producing anything by it's own, hence it consumes 105 chairs.

3): Since, Netherlands gets 15 tables from US and was producing 10 tables from it's own, hence it has now 10+15 = 25 tables.

4); Since, Netherlands gave away 105 chairs of 380 chairs it was producing, hence it has now 380-105 = 275 chairs.